OpenStudy (anonymous):

In order to solve the literal equation A equals s squared for s, what would you do to the equation?

OpenStudy (anonymous):

First set-up your equation\[A = s^2\]

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

Now solve for s, you can do anything to the left side of the equation as long as you also do it to the right side

OpenStudy (anonymous):

Alright, wait. I can do anything?

OpenStudy (anonymous):

Yes but there is something specifically you can do to get the "s" by itself without the 2 as the exponent

OpenStudy (anonymous):

Do you know how to get rid of exponents?

OpenStudy (anonymous):

No, my teacher never taught us. So my mom put me in virtual school so I can start my freshman year on a good hand

OpenStudy (anonymous):

Alright well let us look at the exponent as a fraction\[s^2 = s^\frac{ 2 }{ 1 }\] Does that make sense? since anything over 1 is itself

OpenStudy (anonymous):

To get rid of exponents you must raise both sides of the equation by the reciprocal of the exponent you are trying to get rid of so to review reciprocals: \[\frac{ a }{ b } \] the reciprocal would be the fraction "flipped" like this\[\frac{ b }{ a }\]

OpenStudy (anonymous):

Well, yes. The 2/1 turns into a 2, correct?

OpenStudy (anonymous):

Yes but we can also look at that 2 as 2/1 so that way we can "flip it" and bring it over to the left side

OpenStudy (anonymous):

but an easier way to understand this is that the opposite of exponents is roots

OpenStudy (anonymous):

The root would be ''s''?

OpenStudy (anonymous):

No the root would be the number in the exponent so let me give you an example: \[a = b^4\] To solve for b means that i want to get be on the right side all by itself without the exponent so i must do the opposite of an exponent to get rid of it on the right \[\sqrt[4]{b^4} = b\] but anything i do to the right i must do to the left so \[\sqrt[4]{a} = b\]

OpenStudy (anonymous):

Hmm, okay..

OpenStudy (anonymous):

Using the same idea can you solve for \[A = b^2\]

OpenStudy (anonymous):

\[\sqrt[2]{a}=b\]

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

So the answer is D? My first choice was C...which is ''take the square root of both sides'' D is ''Square both sides''

OpenStudy (anonymous):

when you have more than 1 number for example and they are added or subtract you must do it to both like this \[a + b = c^2\] \[\sqrt[2]{a}+\sqrt[2]{b} = c\] if there are numbers that are multiplied you must put both under root like this\[a \times b + c = d^2\] \[\sqrt[2]{a \times b} + \sqrt[2]{c} = d\]

OpenStudy (anonymous):

No it is was your first choice since you must take the "square root" of "both" sides

OpenStudy (anonymous):

Alright, thank you so much for your help!!!

OpenStudy (anonymous):

square root is 2 outside of the root sign\[\sqrt[2]{a}\] cube root is 3 outside of the root sign \[\sqrt[3]{a}\] and the terms can go on and on